Suppose that GDP is $8 billion, taxes are $1.5 billion, private saving is $0.5 billion, and
public saving is 0.2 billion. Assuming the economy is closed, calculate the size of:


- Disposal income

- Consumption
- Investment
- Government Spending
- National Savings
Answer #1

GDP is 8 billion. Taxes are 1.5 billion. Private saving is 0.5 billion. Public saving is 0.2 billion.

National savings = Private saving + Public saving = 0.5 + 0.2 = 0.7 billion

Government spending = Taxes - Public savings = 1.5 - 0.2 = 1.3 billion

Investment = National savings = 0.7 billion

Disposable income = GDP - taxes = 8 - 1.5 = 6.5 billion

Consumption = Disposable income - Private savings = 6.5 - 0.5 = 6 billion.